# DICE: End-to-end Deformation Capture of Hand-Face Interactions from a Single Image

### v1

@@ 3.2 Weakly-Supervised Training Scheme
Although the aforementioned benchmark, Decaf [97], accurately captures hand, face, self-contact, and deformations, it consists of eight subjects and is recorded in a controlled environment with green screens. Training a model only with Decaf limits its generalization capability to in-the-wild images that have far more complex and diverse human identities, hand poses, and face poses.
To further enhance the generalization capability of our model, we train our model with $500$ diverse in-the-wild images of hand-face interaction collected from the internet without the reliance on the 3D ground truth annotations. First, we use 2D hand-and-face keypoints detected by [68] and [3].
Then, we propose to use Marigold [44], a diffusion-based monocular depth estimator pre-trained on a large number of images to generate 2D affine-invariant depth maps for supervision in the direction of depth (see Eq. 4). The depth supervision provides a strong depth prior, which guides the spatial relationship between hand and face meshes, promoting accurate modeling of hand-face interaction. For supervision, we use a differentiable rasterizer [89] to compute a depth map from the predicted hand and face meshes and supervise the network using a depth loss calculated between the depth values of hand and face keypoints and corresponding points on the predicted depth map. The introduced weak-supervision pipeline significantly enhances our model’s generalization capability and robustness, which we investigate in Sec. 4.4. In our experiment, we found that when training the model with only a small dataset of 500 images, we could significantly improve the model’s accuracy and generalization capability. Moreover, we train adversarial priors on the hand and face parameter space on multiple hand and face pose datasets: the face-only RenderMe-360 [79], the hand-only FreiHand [140], and Decaf [97]. This ensures the plausibility of generated face and hand poses and shapes while allowing for flexible poses and shapes beyond the Decaf data distribution to handle in-the-wild cases.
@@ 3.3 Loss Functions
Mesh losses $\mathcal{L_{\text{mesh}}}$: For richly annotated data in Decaf [97], we employ $L_{1}$ loss for 3D keypoints, 3D vertices, and 2D reprojected keypoints against their respective ground-truths, following common practice in human- and hand-mesh recovery [63, 11, 18]. We further apply a $L_{1}$ loss $\mathcal{L}_{\text{params}}$ on the estimated hand and face pose, shape, and facial expression against the ground-truth parameters. For in-the-wild data, only the 2D reprojected keypoints are supervised, as this is the only type with corresponding ground truth.
Interaction losses $\mathcal{L_{\text{interaction}}}$: For data in Decaf [97], we impose Chamfer Distance losses to enforce touch for predicted contact vertices and discourage collision. We also have a binary cross-entropy loss to supervise contact labels and a deform loss with adaptive weighting to supervise deform vectors, similar to [97]. For in-the-wild data, we also impose touch and collision losses since they do not require annotations.
Adversarial loss $\mathcal{L}_{\text{adv}}$ are applied to the predicted hand and face parameters for in-the-wild data to constrain their parameter space, and for decaf data to facilitate the training of the discriminators. The adversarial loss is given by:
$\mathcal{L}_{\text{adv}}(E)=\mathbb{E}_{\theta_{F}\sim p_{E}}\left[\log\left(1-D_{F}(E(I))\right)\right]+\mathbb{E}_{\theta_{H}\sim p_{E}}\left[\log\left(1-D_{H}(E(I))\right)\right].$ (1)
@@
$\mathcal{L}_{\text{adv}}(D_{F})=-\left(\mathbb{E}_{\theta_{F}\sim p_{E}}\left[\log\left(1-D_{F}(E(I))\right)\right]+\mathbb{E}_{\theta_{F}\sim p_{\text{data}}}\left[\log\left(D_{F}(\theta)\right)\right]\right)$ (2)
@@
$\mathcal{L}_{\text{adv}}(D_{H})=-\left(\mathbb{E}_{\theta_{H}\sim p_{E}}\left[\log\left(1-D_{H}(E(I))\right)\right]+\mathbb{E}_{\theta_{H}\sim p_{\text{data}}}\left[\log\left(D_{H}(\theta)\right)\right]\right),$ (3)
where $E$ jointly denotes the image backbone, the mesh encoder and the parameter regressor, $\theta_{F}=\text{concat}(\theta_{\text{face-shape}},\theta_{\text{face-jaw}},\theta_{\text{face-exp}})$, $\theta_{H}=\text{concat}(\theta_{\text{hand-shape}},\theta_{\text{hand-pose}})$.
Depth loss $\mathcal{L}_{\text{depth}}$: To provide pseudo-3D hand and face keypoints supervision for in-the-wild data, we use a modified SILog Loss [19], an affine-invariant depth loss as our depth supervision $\mathcal{L}_{\text{depth}}$. Formally, let $\hat{K}_{D}$ denote the pseudo-ground-truth affine-invariant depth of the face and hand keypoints, and $K_{D}$ denote the rendered depth for the keypoints,
@@ 4.1 Datasets and Metrics
Datasets
We employ Decaf [97] for reconstructing 3D face and hand interactions with deformations, along with the in-the-wild dataset we collected with 500 images. We use the hand and face shape, pose, and expression data from Decaf [97], RenderMe-360 [79], and FreiHand [140] for training the adversarial priors. We use the training set of the aforementioned datasets for network training. We utilize the Decaf test set for quantitative evaluation, and additionally, we visualize in-the-wild images from the test set for qualitative evaluation.
Metrics.
We adopt commonly-used metrics for mesh recovery accuracy following [43, 63, 18, 11]:
$\bullet$ Mean Per-Joint Position Error (MPJPE): the average Euclidean distance between predicted keypoints and ground-truth keypoints.
$\bullet$ PAMPJPE: MPJPE after Procrustes Analysis (PA) alignment.
$\bullet$ Per Vertex Error: per vertex error (PVE) with translation.
Following Decaf [97], we use the plausibility metrics mentioned below:
$\bullet$ Collision Distance (Col. Dist.): the average collision distances over vertices and frames;
$\bullet$ Non-Collision Ratio (Non. Col.): the proportion of frames without hand-face collisions;
$\bullet$ Touchness Ratio: the ratio of hand-face contacts among ground truth contacting frames;
$\bullet$ F-Score: the harmonic mean of Non-Collision Ratio and Touchness Ratio.
@@ 4.2 Implementation Details
We train the MeshNet, InteractionNet, and IKNet along with the face and hand discriminators with three identical AdamW optimizers with a learning rate of $6\times 10^{-4}$, and a learning rate decay of $1\times 10^{-4}$, optimizing in an alternating manner. Our batch size is set to $16$ during the training stage. Each training task takes $40$ epochs using $48$ hours. The model is trained and evaluated on 8 Nvidia A6000 GPUs with an AMD 128-core CPU. For a fair comparison, all baseline models used the settings in their original papers. Inference times are calculated on a single Nvidia A6000 GPU.
@@ 4.3 Performance on Hand-Face Interaction and Deformation Recovery
In addition to baselines considered in Decaf [97], we compare our method with a representative work in human body/hand mesh recovery, METRO, an end-to-end transformer-based model. For a fair comparison, we compare our method with a modified version of METRO [63] for predicting hand and face meshes, with extra output heads added to predict contact and deformation.
@@
Figure 4: Qualitative comparsion of DICE, Decaf [97], PIXIE [23], METRO*[64] on Decaf validation set and in-the-wild images. Our method achieves superior reconstruction accuracy and plausibility in the Decaf [97] dataset, while generalizing well to difficult in-the-wild actions unseen in Decaf.
@@ 4.3.1 Quantitative Evaluations
Reconstruction Accuracy In Tab. 1, our method surpasses all baseline methods in terms of reconstruction accuracy, achieving a $7.5\%$ reduction in per-vertex error compared to the current state-of-the-art, Decaf. Note that our method is regression-based and allows inference at an interactive rate, while Decaf [97] uses a cumbersome test-time optimization process, taking more than $700$x more time per image. Decaf also requires using temporal information in successive frames, while our method only uses a single frame. Our method shows a $30\%$ reduction in reconstruction error compared to the modified METRO baseline, and up to $79\%$ reduction compared to other end-to-end baselines.
Plausibility In addition to high accuracy, our method achieves the highest overall physical plausibility (F-Score) among all regression-based methods. Note that Touchness and Non-Collision ratio are complement to each other and are meaningless when considered individually, while F-Score measures the two values as a whole. Our method has a much lower interpenetration distance (Col. Dist.) compared to Benchmark and PIXIE (hand+face), which consider hand and face separately, therefore generating implausible interactions. Note that PIXIE (whole body) and METRO* show lower collision distances with a much lower Touchness than our method, indicating that the reconstructed hands and faces often appear incorrectly as if they are not interacting. On the other hand, our method shows low collision distance with a high Touchness, indicating plausible hand-face interaction reconstruction.
Contact Estimation In Tab. 2, DICE achieves superior contact estimation performance on Decaf dataset, surpassing previous work [97] in F-Score for both face and hand contacts. Here, the F-score provides a comprehensive measure of both the precision and the recall ratio combined. These two metrics are complementary and less meaningful when only considered individually; See Fig. 3 for qualitative results.
@@
Decaf (face) [97] $0.57$ $0.69$ $0.49$ $0.99$
Decaf (hand) [97] $0.47$ $0.62$ $0.39$ $0.98$
@@ 4.3.2 Qualitative Evaluations
As discussed in Sec. 3.2, the Decaf [97] dataset is collected in an indoor environment with a green screen, which doesn’t reflect the complex environment where real-world hand-face interactions occur. Therefore, a model only trained with the Decaf dataset might have generalization issues when tested on in-the-wild data. Fig. 4 confirms this result by our model’s superior generalization performance on in-the-wild data with unseen identity and pose. As shown in Fig. 3, our method faithfully reconstructs hand-face interaction and deformation and accurately labels the area of contact.
@@ 4.4 Ablation Study
In-the-wild data As shown in Tab. 3, adding weak-supervision training and in-the-wild data for DICE training improves all reconstruction error metrics (PVE*, MPJPE, PAMPJPE) while maintaining a high plausibility (F-Score). This is because the limited pose and identity distribution of the Decaf training dataset may cause the model to overfit, and the inclusion of in-the-wild images out of the Decaf data distribution effectively improves the generalization capability of DICE.
Depth Supervision Although depth supervision is only applied to in-the-wild data, as shown in Tab. 3, removing it also significantly affects performance on the Decaf validation set. Without depth loss, wrong predictions in depth are not penalized for in-the-wild data, introducing noise in the training process, and resulting in erroneous depth predictions in the Decaf dataset. As shown in Appendix Fig. 7, the absence of depth supervision introduces ambiguity in the z-direction, resulting in artifacts such as self-collision.
Adversarial Prior The adversarial prior incorporates diverse but realistic pose and shape distribution beyond Decaf [97], ensuring the reality of regressed mesh while allowing for generalization. As shown in Tab. 3, introducing adversarial supervision improves the accuracy and physical plausibility.
Parameter Supervision Supervising parameters directly, in addition to the indirect supervision of parameters by the mesh losses, improves both plausibility and accuracy. This is because direct parameter supervision eliminates ambiguity, without which the network may resort to other parameter combinations that produce incorrect meshes similar in terms of Euclidean distance.
Intermediate Supervision Removing the constraint that $\mathbf{V}_{F}^{\prime},\mathbf{V}_{H}^{\prime}$ being rough meshes and treating them only as feature maps results in a substantial drop in accuracy and a slight drop in plausibility (F-Score). Also, note that there is a sharp increase in collision distance, which is attributed to the spatial inaccuracy of the final output mesh. This indicates that using an intermediate mesh feature instead of an ordinary feature map increases the spatial accuracy of the output meshes, which also benefits plausibility.
Network Design In Tab. 3, adopting the two-branch architecture, which separates deformation and interaction estimation from mesh vertices regression, improves both accuracy and plausibility.

### v2

@@ 3.2 Weakly-Supervised Training Scheme
Although the aforementioned benchmark, Decaf (Shimada et al., 2023), accurately captures hand, face, self-contact, and deformations, it consists of only eight subjects and is recorded in a green-screen studio. Thus, training a model only with the Decaf dataset limits its generalization capability to in-the-wild images that exhibit far more complex and diverse human identities, hand poses, and face poses.
To further enhance the generalization capability, we train our model with $500$ diverse in-the-wild images of hand-face interaction collected from the internet without the reliance on the 3D ground truth annotations. First, we use 2D hand and face keypoints detected by Lugaresi et al. (2019) and Bulat & Tzimiropoulos (2017) as pseudo-ground-truth. Then, we use Marigold (Ke et al., 2024), a diffusion-based monocular depth estimator pre-trained on a large amount of images to generate 2D affine-invariant depth maps for depth supervision (see Eq. 4). The depth supervision provides a strong depth prior, which guides the spatial relationship between hand and face meshes, promoting accurate modeling of hand-face interaction. We first use a differentiable rasterizer (Ravi et al., 2020) to compute a depth map from the predicted hand and face meshes. We use a depth loss to measure the difference between the depths of the hand and face keypoints and their corresponding points on the predicted depth map, providing supervision. This keypoint-to-keypoint correspondence enables accurate depth supervision even when the rendered hand/face mesh and the ground-truth meshes are misaligned. Moreover, we train adversarial priors on the hand and face parameter space on multiple hand and face motion datasets: the face-only RenderMe-360 (Pan et al., 2023a), the hand-only FreiHand (Zimmermann et al., 2019), and Decaf (Shimada et al., 2023). This ensures the plausibility of generated face and hand poses and shapes while allowing for flexible poses and shapes beyond the Decaf data distribution to handle in-the-wild cases. The overall weak-supervision pipeline significantly enhances our model’s generalization capability and robustness, which we investigate in Sec. 4.4.
@@ 3.3 Loss Functions
Mesh losses $\mathcal{L_{\text{mesh}}}$: For richly annotated Decaf dataset (Shimada et al., 2023), we employ an $L_{1}$ loss for 3D keypoints, 3D vertices, and 2D reprojected keypoints, comparing them against their respective ground-truths, following common practice in human and hand mesh recovery (Lin et al., 2021a; Cho et al., 2022; Dou et al., 2023b). We further apply an $L_{1}$ loss $\mathcal{L}_{\text{params}}$ on the estimated hand and face pose, shape, and facial expression against the ground-truth parameters. For in-the-wild data, only the 2D reprojected keypoints are supervised, as they are the only type with corresponding ground truth.
Interaction losses $\mathcal{L_{\text{interaction}}}$: Similar to Shimada et al. (2023), we impose Chamfer Distance losses to promote touch for predicted contact vertices and discourage collision. We also introduce a binary cross-entropy loss to supervise contact labels and a deform loss with adaptive weighting mechanism to supervise deform vectors. For in-the-wild data, we also impose touch and collision losses since they do not require annotations.
Adversarial loss $\mathcal{L}_{\text{adv}}$: The adversarial losses are applied to the predicted hand and face parameters for in-the-wild data to constrain their parameter space, and for Decaf data to facilitate the training of the discriminators. The adversarial loss is given by:
$\mathcal{L}_{\text{adv}}(E)=\mathbb{E}_{\bm{\theta}_{f}\sim p_{E}}\left[\log\left(1-D_{F}(E(I))\right)\right]+\mathbb{E}_{\bm{\theta}_{h}\sim p_{E}}\left[\log\left(1-D_{H}(E(I))\right)\right].$ (1)
@@
$\mathcal{L}_{\text{adv}}(D_{F})=-\left(\mathbb{E}_{\bm{\theta}_{f}\sim p_{E}}\left[\log\left(1-D_{F}(E(I))\right)\right]+\mathbb{E}_{\bm{\theta}_{f}\sim p_{\text{data}}}\left[\log\left(D_{F}(\bm{\theta}_{f})\right)\right]\right),$ (2)
@@
$\mathcal{L}_{\text{adv}}(D_{H})=-\left(\mathbb{E}_{\bm{\theta}_{H}\sim p_{E}}\left[\log\left(1-D_{H}(E(I))\right)\right]+\mathbb{E}_{\bm{\theta}_{H}\sim p_{\text{data}}}\left[\log\left(D_{H}(\bm{\theta}_{h})\right)\right]\right),$ (3)
where $E$ jointly denotes the image backbone, the mesh encoder and the parameter regressor, $p_{E}$ denotes the output distribution of $E$, $p_{\text{data}}$ denotes the data distribution of the motion datasets, $\bm{\theta}_{f}=(\bm{\theta}_{\text{f-pose}},\bm{\beta}_{\text{f}},\bm{\theta}_{\text{f-exp}})$, $\bm{\theta}_{H}=(\bm{\theta}_{\text{h}},\bm{\beta}_{\text{h}})$.
Depth loss $\mathcal{L}_{\text{depth}}$: To provide pseudo-3D hand and face keypoints supervision for in-the-wild data, we use a modified SILog Loss (Eigen et al., 2014), an affine-invariant depth loss as our depth supervision $\mathcal{L}_{\text{depth}}$. Formally, let $\hat{K}_{D}$ denote the pseudo-ground-truth affine-invariant depth of the face and hand keypoints, and $K_{D}$ denote the rendered depth for the keypoints,
@@ 4.1 Datasets and Metrics
Datasets. We employ Decaf (Shimada et al., 2023) for reconstructing 3D face and hand interactions with deformations, along with the in-the-wild dataset we collected containing 500 images. We use the shape, pose, and expression data of hands and faces from Decaf (Shimada et al., 2023), RenderMe-360 (Pan et al., 2023a), and FreiHand (Zimmermann et al., 2019) for training the adversarial priors. We use the training set of the aforementioned datasets for network training. We use the official split from Decaf to separate the training and testing sets, and select a few in-the-wild images for the test set to perform qualitative visualizations.
Metrics. We adopt commonly-used metrics for mesh recovery accuracy following Kanazawa et al. (2018); Lin et al. (2021a); Dou et al. (2023b); Cho et al. (2022): 1) Mean Per-Joint Position Error (MPJPE): the average Euclidean distance between predicted keypoints and ground-truth keypoints. 2) PAMPJPE: MPJPE after Procrustes Analysis (PA) alignment. 3) Per Vertex Error: per vertex error (PVE) with translation. Following Decaf (Shimada et al., 2023), we use the following metrics to measure the plausibility: 4) Collision Distance (Col. Dist.): the average collision distances over vertices and frames; 5) Non-Collision Ratio (Non. Col.): the proportion of frames without hand-face collisions; 6) Touchness Ratio (Touchness): the ratio of hand-face contacts among ground truth contacting frames; 7) F-Score: the harmonic mean of Non-Collision Ratio and Touchness Ratio. Note that F-Score measures Touchness and Non-Collision Ratio as a whole, which is a metric of overall physical plausibility, whereas Non-Collision Ratio or Touchness are meaningless when considered individually.
@@ 4.2 Implementation Details
We train MeshNet, InteractionNet, and IKNet, along with the face and hand discriminators using AdamW (Loshchilov, 2017) optimizers, each with a learning rate of $6\times 10^{-4}$, and a learning rate decay of $1\times 10^{-4}$. The generator and discriminator networks are optimized in an alternating manner. Our batch size is set to $16$ during the training stage. The training takes $40$ epochs, totalling $48$ hours. The model is trained and evaluated on 8 Nvidia A6000 GPUs with an AMD 128-core CPU. Inference times are calculated on a single Nvidia A6000 GPU.
@@ 4.3 Performance on Hand-Face Interaction and Deformation Recovery
We compare our method with the following: 1) Benchmark: the baseline (Lugaresi et al., 2019; Li et al., 2017) introduced in Decaf (Shimada et al., 2023); 2) Decaf (Shimada et al., 2023): an optimization-based method for hand-face interaction and deformation recovery. 3) PIXIE (whole-body) (Feng et al., 2021a): a representative model for full-body recovery, including the hand and face, introduced in Decaf. 4) PIXIE (hand+face) (Feng et al., 2021a): a optimization-based variant of PIXIE, introduced in Decaf. For regression-based methods, as we are dealing with a relatively new task, there are few readily available baselines. To facilitate comparison, we adapt the following regression-based models from related tasks: 5) METRO (Lin et al., 2021a): A representative work in human body/hand mesh recovery. We adapt METRO to predict both hand and face meshes, adding extra output heads to predict contact and deformation. 6) PIXIE-R (Feng et al., 2021a): Adapted PIXIE, using the same backbone and hand and face branches but trained with losses from DICE. 7) FastMETRO (single-target) (Cho et al., 2022): Another representative work in human and hand mesh recovery. We adapt two independent FastMETROs, one for estimating hand mesh vertices and contact, and the other for estimating face mesh, deformation, and contact. Here, the word single-target means each FastMETRO considers hand and face individually, with no information exchange. This model is trained using the same hyperparameter, loss, and optimizer as DICE, on the Decaf (Shimada et al., 2023) dataset.
@@
Figure 4: Qualitative comparison of DICE, Decaf (Shimada et al., 2023), PIXIE (Feng et al., 2021a) (whole-body version), METRO* (Lin et al., 2021b) on Decaf validation set and in-the-wild images. Our method achieves superior reconstruction accuracy and plausibility in the Decaf (Shimada et al., 2023) dataset, especially generalizing well to difficult in-the-wild actions unseen in Decaf compared to all baselines.
@@ 4.3.1 Quantitative Evaluations
Reconstruction Accuracy. In Tab. 1, our method surpasses all baseline methods in terms of reconstruction accuracy, achieving a $7.5\%$ reduction in per-vertex error compared to the current state-of-the-art, Decaf. Note that our method is regression-based and allows inference at an interactive rate, while Decaf (Shimada et al., 2023) uses a cumbersome test-time optimization process, taking more than $200$x more time per image. Decaf also requires using temporal information in successive frames, while our method only uses a single frame. Our method shows a $30\%$ reduction in reconstruction error compared to the modified METRO baseline, and up to $79\%$ reduction compared to other end-to-end baselines. Notably, our method achieves a $27\%$ MPVE reduction compared to the PIXIE-R baseline which uses the same mesh and interaction losses as our method, demonstrating the superiority of our network design and weak-supervised training scheme. Our method is also more accurate than another end-to-end baseline, FastMETRO.
Plausibility. In terms of overall physical plausibility (F-Score), our method is the best among all regression-based methods: PIXIE (whole-body), PIXIE-R, METRO, and FastMETRO. On the other hand, while some optimization-based methods (Decaf and PIXIE (hand+face) have higher overall plausibility (F-Score) compared to DICE, this is due to their test-time optimization, which iteratively adjusts the relative positioning of hand and face. Thus, they are much more computationally intensive than our regression-based method. With a highly efficient end-to-end inference scheme, DICE still outperformed an optimization-based method (Benchmark) on F-Score.
Contact Estimation. The contact estimation metrics (accuracy, precision, recall) are calculated by the predicted per-vertex contact probabilities against the respective 0-1 contact ground truths. In Tab. 2, DICE achieves superior contact estimation performance on the Decaf dataset, surpassing previous work (Shimada et al., 2023) in F-Score for both face and hand contacts. Here F-score provides a comprehensive measure of both precision and recall ratio combined. These two metrics involve a trade-off: focusing solely on precision may lead to a decrease in recall, and vice versa. Balancing this trade-off, the F-score offers a more meaningful evaluation of contact estimation.
Table 2: Comparison of hand-face contact estimation on Decaf.
@@
Decaf (face) (Shimada et al., 2023) $0.57$ $0.69$ $0.49$ $0.99$
Decaf (hand) (Shimada et al., 2023) $0.47$ $0.62$ $0.39$ $0.98$
@@ 4.3.2 Qualitative Evaluations
As discussed in Sec. 3.2, the Decaf (Shimada et al., 2023) dataset is collected in an indoor environment with a green screen, which doesn’t reflect the complex environment where real-world hand-face interactions occur. Therefore, a model only trained with the Decaf dataset might have generalization issues when tested on in-the-wild data. Fig. 4 supports this claim by demonstrating our model’s superior generalization performance on in-the-wild data with unseen identity and pose. On the other hand, Decaf’s reconstruction suffers from self-collision and incorrect hand-face relationship. PIXIE and METRO reconstruct inaccurate hand poses and often demonstrates implausible non-touching artefacts. As shown in Fig. 3, our method faithfully reconstructs hand-face interaction and deformation and accurately labels the contact areas.
@@ 4.4 Ablation Study
Network Design. In Tab. 3, adopting the two-branch architecture, which separates deformation and interaction estimation from mesh vertices regression, improves both accuracy and plausibility.
In-the-wild data. As shown in Tab. 3, adding weak-supervision training and in-the-wild data for DICE training improves all reconstruction error metrics (PVE*, MPJPE, PAMPJPE) while maintaining a high plausibility (F-Score). We deem that the slight decrease in F-Score could mainly be attributed to the difference in distribution between the studio-collected Decaf (Shimada et al., 2023) and in-the-wild data. This is because the limited pose and identity distribution of the Decaf training dataset may cause the model to overfit, and the inclusion of in-the-wild images out of the Decaf data distribution effectively improves the generalization capability of DICE.
Unrefined Features Supervision. Regressing the unrefined head and hand mesh features $\mathbf{V}_{F}^{\prime},\mathbf{V}_{H}^{\prime}$ and then perform inverse kinematics to regress the parametric mesh improves plausibility and accuracy, compared to directly estimating the face and hand parameters.
Depth Supervision. Although depth supervision is only applied to in-the-wild data, as shown in Tab. 3, removing it also significantly degrades performance on the Decaf validation set. Without depth loss, wrong predictions in depth are not penalized for in-the-wild data, introducing noise in the training process, and resulting in erroneous predictions on the Decaf dataset. As shown in Appendix Fig. 7, the absence of depth supervision introduces ambiguity in the z-direction, resulting in artifacts such as self-collision.
Parameter Supervision. Supervising parameters directly, in addition to the indirect supervision of parameters by the mesh losses, improves both plausibility and accuracy. This is because direct parameter supervision eliminates ambiguity, preventing the network from converging to alternative parameter combinations that produce incorrect meshes that appear geometrically similar, i.e., with small vertex loss, to the target but are incorrect in their underlying structure, such as pose or shape.
Adversarial Prior. The adversarial prior incorporates diverse but realistic pose and shape distribution beyond Decaf (Shimada et al., 2023), ensuring the reality of regressed mesh while allowing for generalization. As shown in Tab. 3, introducing adversarial supervision improves the accuracy and physical plausibility.